# Integrated Sensing and Communications Over the Years: An Evolution Perspective

### v2

@@ VI-A2 Brief Review of IEEE Standards
In parallel with 3GPP, IEEE has made significant strides in ISAC standardization. A major milestone is the IEEE 802.11bf standard, which incorporates sensing functionalities into the existing Wi-Fi ecosystem and contributes to the broader ISAC framework. Managed by the IEEE 802.11 task group bf, this effort aims to formalize Wi-Fi-based sensing within the IEEE 802.11 family, complementing 3GPP’s efforts [233].
@@
In 2018, IEEE began exploring Wi-Fi sensing by leveraging CSI and the received signal strength indicator (RSSI) for environmental detection. Early studies demonstrated the feasibility of using reflected Wi-Fi signals to detect motion, presence, and physiological parameters without the need for additional hardware. Key use cases included indoor localization, smart home applications, healthcare monitoring, and gesture recognition. To formalize these efforts, the IEEE 802.11bf was established in 2020 to standardize Wi-Fi sensing. The initiative aimed to integrate sensing functionalities into existing Wi-Fi standards while ensuring backward compatibility, minimizing resource overhead to preserve communication performance, and optimizing ambiguity functions to enhance sensing accuracy through waveform design and reflection analysis.
@@
By 2022, the IEEE 802.11bf released draft standards focusing on several key aspects:
  • Channel Measurement Support: The use of existing Wi-Fi signals, such as training fields in preambles and pilot tones, was standardized to enable sensing without altering the fundamental Wi-Fi frame structure. The introduction of sensing measurement reports facilitated the exchange of channel measurements for environmental interpretation.
  • Waveform Optimization: OFDM-based Wi-Fi waveforms were refined to improve sensing accuracy without compromising data rates. Improvements in ambiguity functions enabled better separation of multiple sensing objects.
  • Interference Mitigation: Mechanisms were developed to balance S&C by optimizing resource allocation. Dynamic allocation of spectrum and time resources was proposed to support simultaneous sensing and data transmission.
@@
Currently, the IEEE 802.11bf is refining the draft standards, with final approval anticipated in 2024. Efforts are concentrated on enhancing channel measurement techniques for precise location-based services and aligning with adjacent fields such as autonomous systems and extended reality. In addition, multi-user resource management is being optimized to support dense deployment scenarios, which are critical for robust and scalable ISAC applications.
@@ VI-A3 Brief Review of ITU Standards
At the World Radiocommunication Conference 2019 (WRC-19), ITU allocated higher frequency bands, including mmWave spectrum in the 24–52.6 GHz range, to support 5G. These allocations are highly favorable for ISAC, as the wide bandwidth and short wavelengths enable fine sensing resolution. In the next phase, ITU has also initiated studies on sub-terahertz (sub-THz) and terahertz (THz) bands (100–300 GHz), which are expected to be critical for high-resolution S&C in 6G networks [234].
@@
ITU’s contributions center on spectrum allocation, global harmonization, and aligning standards with ISAC objectives. Prioritizing smart spectrum coexistence and dynamic sharing, ITU advanced frameworks support ISAC integration. Furthermore, ITU has initiated discussions on the role of metasurfaces and RIS, highlighting their potential to enhance sensing resolution and optimize the performance of ISAC systems.
@@ VI-B Lessons Learned
Standardization has advanced ISAC from a conceptual vision to a pre-commercial reality, yet several technical and organizational challenges remain. Insights from current standardization efforts, aligned with emerging research demands, point toward the following key questions:
@@ VI-B1 How Can ISAC Performance be Validated through Standardized KPIs
Standardization bodies have defined performance metrics and testing methodologies to enable systematic evaluation in ISAC. IEEE 802.11bf specifies sensing KPIs such as velocity (0.1–0.4 m/s) and angular resolution (1°–8°), which are critical for assessing the trade-offs between S&C performance. 3GPP further incorporates these KPIs into channel models to validate ISAC under real-world cellular environments, particularly in vehicular networks and IoT systems [235].
@@ VI-B2 How Can Use Case-Driven Research Guide Standardization
3GPP SA1 has consolidated extensive research on ISAC scenarios within the RAN1 channel modeling project, with industry consensus highlighting the pivotal roles of the low-altitude economy, drones, and intelligent transportation. For example, enhancements in 5G NR Release 18 enable drones to support beamforming, allowing BSs to detect unauthorized flights, perform obstacle avoidance, and guide trajectories via multi-station collaboration. Integrating passive tracking with active sensing improves positioning accuracy, while environmental awareness assists beam management and mobility. Furthermore, mmWave operation achieves low latency with slot lengths as short as 0.125 ms (per 3GPP TS 38.211), ensuring real-time responsiveness.
The low-altitude economy has also gained strong policy support. In China, the Civil Aviation Administration has promoted 5G NR–based applications, and the Ministry of Industry and Information Technology announced on October 23 initiatives to establish ISAC-enabled infrastructure for low-altitude networks [^1] https://english.www.gov.cn/news/202410/23/content_WS6718e264c6d0868f4e8ec381.html. Despite these advances, practical deployments face challenges such as dense connectivity, interference, and spectrum conflicts. To address these, joint optimization methods—such as beamforming, dynamic spectrum allocation, and coordinated frequency management—have been explored. For instance, [236] proposed a joint UAV maneuver and beamforming strategy that improved trajectory planning, bandwidth allocation, and overall energy efficiency.
@@ VII Conclusion and Future Directions
This survey has extensively examined the impact of ISAC from multiple dimensions, including spectrum utilization, network architectures, sensing methodologies, security concerns, and ongoing standardization efforts. Collectively, these developments illustrate both the transformative potential of ISAC and the opportunities it offers across multiple applications.
@@
As ISAC matures, its role in wireless communication will likely expand, shaping more efficient, intelligent, and adaptive networks. By embedding sensing functions into communication infrastructures, ISAC offers the foundation for wireless systems that actively interpret and respond to their environments in real time. Nevertheless, this evolution introduces substantial challenges, including sophisticated resource management, robust security mechanisms, and scalable hardware integration.
@@
  • Unified channel modeling: Unlike traditional communication models, ISAC requires hybrid geometric–stochastic representations that account for multipath reflections, clutter, and object-induced scattering. While the ongoing 3GPP RAN1 study [230] marks an important first step, future standards must adopt multi-bounce, clutter-aware, and object-tagged models to support sub-THz and mmWave sensing with high fidelity.
  • Resource optimization under multi-objective constraints: Emerging techniques such as rate-splitting multiple access (RSMA) [237] and OTFS waveform design [238] offer promising solutions, yet current scheduling frameworks rarely consider sensing-specific KPIs, such as angular resolution or Doppler sensitivity. Developing ISAC-aware scheduling policies that dynamically balance spatial, temporal, and spectral resources represents a compelling direction for ongoing research.
  • Privacy and Security: The exposure of environmental and user-related data introduces significant risks. Beyond conventional encryption, sensing-aware systems require native security features, such as federated access control, encrypted beamforming, and zero-trust architectures [239]. At the same time, these mechanisms must align with evolving global data protection regulations to ensure lawful and ethical deployment.

### v3

@@ VI-A2 Brief Review of IEEE Standards
    In parallel with 3GPP, IEEE has made significant strides in ISAC standardization. A major milestone is the IEEE 802.11bf standard, which incorporates sensing functionalities into the existing Wi-Fi ecosystem and contributes to the broader ISAC framework. Managed by the IEEE 802.11 Task Group bf, this effort aims to formalize Wi-Fi-based sensing within the IEEE 802.11 family, complementing 3GPP’s efforts [kaushik2024toward].
@@
    In 2018, IEEE began exploring Wi-Fi sensing by leveraging CSI and the received signal strength indicator (RSSI) for environmental detection. Early studies demonstrated the feasibility of using reflected Wi-Fi signals to detect motion, presence, and physiological parameters without the need for additional hardware. Key use cases included indoor localization, smart home technologies, healthcare surveillance, and gesture recognition. The IEEE 802.11bf was instituted in 2020 to standardize Wi-Fi sensing. The initiative aimed to integrate sensing capabilities into existing Wi-Fi standards while ensuring backward compatibility and minimizing resource overhead so that communication performance is preserved. It also improved the ambiguity function through optimized waveform design and reflection analysis, leading to enhanced sensing accuracy.
@@
    By 2022, the IEEE 802.11bf draft standard had begun defining several foundational elements for Wi-Fi sensing. It formalized the use of existing Wi-Fi waveform components, including preamble training fields and pilot tones, to enable sensing without modifying the underlying frame structure, and it introduced sensing measurement reports to support the exchange of channel information for environmental inference. Refinements to OFDM-based waveforms improved sensing accuracy while preserving throughput, and enhancements to the ambiguity function enabled more reliable multi-target separation. The draft proposed mechanisms for interference mitigation, outlining dynamic strategies for allocating spectrum and time resources so that sensing operations and data transmission could coexist efficiently.
@@
    The IEEE 802.11bf is refining the draft standards, with final approval anticipated in 2024. Efforts are concentrated on enhancing channel measurement techniques for precise location-based services and on aligning with adjacent fields such as autonomous systems and extended reality. In addition, multi-user resource management is being optimized to support dense deployment scenarios, which are critical for robust and scalable ISAC applications.
@@ VI-A3 Brief Review of ITU Standards
    At the World Radiocommunication Conference 2019 (WRC-19), ITU allocated higher frequency bands, including mmWave spectrum in the 24–52.6 GHz range, to support 5G. These allocations are highly favorable for ISAC, as the wide bandwidth and short wavelengths enable fine sensing resolution. In the next phase, ITU has also initiated studies on sub-terahertz (sub-THz) and terahertz (THz) bands (100-300 GHz), which are expected to be critical for high-resolution S&C in 6G networks [kaushik2024toward].
@@
    ITU’s contributions center on spectrum allocation, global harmonization, and aligning standards with ISAC objectives. Prioritizing smart spectrum coexistence and dynamic sharing, ITU advanced frameworks support ISAC integration. ITU has also initiated discussions on the role of metasurfaces and RIS, highlighting their potential to enhance sensing resolution and optimize the performance of ISAC systems.
@@ VI-B Lessons Learned
    Standardization has advanced ISAC from a conceptual vision to a pre-commercial reality, yet several technical and organizational challenges remain. Insights from current standardization efforts, aligned with emerging research demands, point to the following key questions:
@@ VI-B1 How Can ISAC Performance be Validated through Standardized KPIs
    Standardization bodies have defined performance metrics and testing methodologies to enable systematic evaluation of ISAC. IEEE 802.11bf specifies sensing KPIs such as velocity (0.1–0.4 m/s) and angular resolution (1°–8°), which are critical for assessing the trade-offs between S&C performance. 3GPP further incorporates these KPIs into channel models to validate ISAC under real-world cellular environments, particularly in vehicular networks and IoT systems [10251772]. These benchmarks ensure that standardized ISAC solutions achieve acceptable trade-offs between sensing accuracy and communication reliability.
    Figure 14: ISAC-based rotary-wing UAV Scenario [10529184]
@@ VI-B2 How Can Use Case-Driven Research Guide Standardization
    3GPP SA1 has consolidated extensive research on ISAC scenarios within the RAN1 channel-modeling project, with industry consensus highlighting the pivotal roles of the low-altitude economy, drones, and intelligent transportation. For example, enhancements in 5G NR Rel-18 enable drones to support beamforming, allowing BSs to detect unauthorized flights, perform obstacle avoidance, and guide trajectories via multi-station collaboration. Integrating passive tracking with active sensing improves positioning accuracy, while environmental awareness assists beam management and mobility. Furthermore, mmWave operation achieves low latency with slot lengths as short as 0.125 ms, ensuring real-time responsiveness.
    The low-altitude economy has also gained strong policy support. In China, the Civil Aviation Administration has promoted 5G NR–based applications, and the Ministry of Industry and Information Technology announced on October 23 initiatives to establish ISAC-enabled infrastructure for low-altitude networks [^1] https://english.www.gov.cn/news/202410/23/content_WS6718e264c6d0868f4e8ec381.html. Despite these advances, practical deployments face challenges such as dense connectivity, interference, and spectrum conflicts. To address these, joint optimization methods have been explored, including beamforming, dynamic spectrum allocation, and coordinated frequency management. For instance, [lyu2022joint] proposed a joint UAV maneuver and beamforming strategy that improved trajectory planning, bandwidth allocation, and overall energy efficiency. These optimization techniques significantly enhance system performance, as demonstrated in [10529184] and Fig. 14.
@@ VII Conclusion and Future Directions
    This paper has extensively examined the impact of ISAC across key domains, including operating spectrum, network architecture, sensing methodologies, security challenges, and standardization efforts within the context of 5G and beyond. These developments provide insights into ISAC’s potential and the opportunities it offers across multiple applications.
@@
    As ISAC matures, its role in wireless communication will likely expand, shaping more efficient, intelligent, and adaptive networks. By embedding sensing functions into the communication infrastructure, ISAC provides wireless systems with the ability to understand and respond to the environment in real time. However, this evolution has also brought many challenges, including complex resource management mechanisms, higher-intensity security requirements, and scalable hardware integration schemes.
@@
    • Unified channel modeling: Unlike traditional communication models, ISAC demands hybrid geometric-stochastic representations that take into account multipath reflections, clutter, and scattering caused by objects. Although the continuous research of 3GPP RAN1 [R1-240] has taken an important step, future standards still need to adopt multi-bounce, clutter aware and object-tagged models to support sub-THz and mmWave sensing with high fidelity.
    • Resource optimization under multi-objective constraints: Emerging technologies such as rate-splitting multiple access (RSMA) [xu2021rate] and OTFS waveform design [gaudio2020effectiveness] show good prospects, but the existing scheduling framework rarely incorporates KPIs specific to sensing, such as angle resolution or Doppler sensitivity. Developing ISAC-aware scheduling policies that dynamically balance spatial, temporal, and spectral resources represents a compelling direction for ongoing research.
    • Privacy and Security: The exposure of environmental and user-related data introduces significant risks. Beyond conventional encryption, sensing-aware systems require native security features, such as federated access control, encrypted beamforming, and zero-trust architectures [wei2022toward]. At the same time, these mechanisms must align with evolving global data protection regulations to ensure lawful and ethical deployment.